4. Cost of bond. XYZ Corporation has an outstanding 10-year bond with a 6% coupon rate which sells for $1,015. The company has a 35% tax rate. How much does the bond cost to the company? 3.77% With detailed calculated steps.
Answer #1
                  K = N
Bond Price =∑ [(Annual Coupon)/(1 + YTM)^k]     +   Par value/(1 + YTM)^N
                   k=1
                  K =10
1015 =∑ [(6*1000/100)/(1 + YTM/100)^k]     +   1000/(1 + YTM/100)^10
                   k=1
YTM% = 5.8
Using Calculator: press buttons "2ND"+"FV" then assign
PV =-1015
PMT = Par value * coupon %=1000*6/(100)
N =10
FV =1000
CPT I/Y
Using Excel
=RATE(nper,pmt,pv,fv,type,guess)
=RATE(10,-6*1000/(100),1015,-1000,,)
After tax rate = YTM * (1-Tax rate)
After tax rate = 5.8 * (1-0.35)
After tax rate = 3.77